OpenStudy (anonymous):

-5x-5=? show your work

OpenStudy (anonymous):

\[-5x=5\] \[x = -5/5\] x= -1

OpenStudy (anonymous):

is that supposed to say equal to 0? if so then perform the inverse operations. add 5 to both sides then divide by -5 to get x

OpenStudy (anonymous):

hurry i have one medal

OpenStudy (anonymous):

Is that the full equation? If so, -5x-5=0 add 5 to both sides; -5x=5 divide both sides by -5 x = -1

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

all you wrong

OpenStudy (anonymous):

Did you enter the right equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well since ? isn't a quantity we will never know what your trying to equal

OpenStudy (anonymous):

try again! -5x-5= -60 show work

OpenStudy (anonymous):

well then looks like its going to be 11

OpenStudy (anonymous):

-5x - 5 = -60 add 5 to both sides -5x = -55 divide both sides by -5 x = 11

OpenStudy (anonymous):

you winning ryan!